OpenStudy (anonymous):

substitution method: 2x-y=4, x+2y=7

OpenStudy (anonymous):

Ok. First 2x=y+4 then x= (y+4)/2 from this point ; substitute this value to the second question ; (y+4)/2 +2y =7 then equalise the denominators. it will be like (5y + 4 ) / 2 = 7 then solve y . which is 14 = 5y + 4 10 = 5y y = 2 then substitute y into the equations like x +2y = 7 then x + 4 = 7 x= 3 thats all :D